OpenStudy (anonymous):

Factor the trinomial 4x2 + 24x + 11 once i get it broken down to 4x^2+22x+2x+11 i am stuck, what is the next step?

OpenStudy (anonymous):

4x^2+24x+11

OpenStudy (anonymous):

(2x+11)(2x +1)

OpenStudy (anonymous):

can you show me how you got that?

OpenStudy (anonymous):

well its trial and error first assume its 2x and 2x in the firstbpart of the bracket now u nedd to numbers which when multiplied together give 11, and when multiplied by 2 and added give 24 these numbers are 1 and 11 OUtside numbers give 2x * 11 = 22x and inside give 2x * 1 = 2x 2x + 22x = 24x

OpenStudy (anonymous):

i was just confused becaquse my book told me to multiple 4x and 11 as the first step so i got 44x which would be x+22 and x+2 cus they add to 24 and mulitply to 44 but then once i split the middle term im stuck

OpenStudy (anonymous):

Ok lets try factoring x^2 - x - 6 first part of brackets will be x (to give x^2 (x )(x ) Now we need 2 numbers which when multip;ies will give -6 and when added will give -1. these 2 numbers are -3 and 2:- -3 * +2 = --6 and -3 + 2 = -1 so our answer is (x -3)(x + 2)

OpenStudy (anonymous):

right i understand that i just dpnt get what to do when that x has a leading coeeficiant like 4x^2+24x+11

OpenStudy (anonymous):

yeah there are a few ways do do this - I personally prefer the way ive shown

OpenStudy (anonymous):

which was the trial and error method?

OpenStudy (amistre64):

step1; factor out commons; nothing factors move to step 2 step2; setup as if it factors: (x ) (x ) step3; multiply the first and last numbers to get a pool of options; and divide out the first from the set up: (x /4) (x /4) 4(11) = 44 step4; observe the signs of the middle and last numbers to fill in the setup; +24; +11. the last number tells us to add; the middle number says the bigger number is +; which would make better sense in another problem, but lets go with it :) (x+B/4) (x+ S/4) step5; go thru the factors to determine the numbers to fill in: 1,44; 1+44 = 45 .... toss it 2,22; 2+22 = 24 .... yay!! we can use it step6 ; fill in the setup (x+ 22/4) (x+ 2/4) step7; reduce till you cant reduce no more (x+ 11/2) (x+1/2) step8; if theres still a fraction; throw the bottom in front (2x +11) (2x +1) step9 ; dont forget to include the factor from step 1; in this case, there is none

OpenStudy (anonymous):

i kind of see it there, im going to try the next problem

OpenStudy (anonymous):

Factor the trinomial 9x2 - 18xy + 5y2

OpenStudy (anonymous):

the one i've demonstrated above. u can write them in columns if u like 4x^2 11 24 2 X 2 11 X 1 2X11 + 2 X 1 = 24

OpenStudy (amistre64):

my setup is: (x y/9)(x y/9) the last sign says + and the middle says - is bigger (x- By/9)(x- Sy/9) 9(5) = 45 1,45; 3,15 ... 3+15 - 18 (x -15y/9)(x -3y/9) (x -5y/3)(x -y/3) (3x -5y) (3x -y)

OpenStudy (anonymous):

your dividing them?

OpenStudy (amistre64):

my way; you multiply the first into the last; so in the end you have to divide it out again

OpenStudy (anonymous):

oh my i am so confused lol

OpenStudy (anonymous):

9 +5 -18 3 X 3 -5 X -1 3 X -5 + 3 X -1 = -18 so the factors are (3x -5y)(3x - y)

OpenStudy (amistre64):

when there is no leading coeff; or rather, it is 1; then you simply address the last number to get your options

OpenStudy (anonymous):

i See it !!! thank you

OpenStudy (amistre64):

this way; you use it to make life easier to begin with; and remember to take it back out when your finished

OpenStudy (anonymous):

i like how its broken down into columns that makes it a lot easier to see

OpenStudy (anonymous):

thats quite a good way amis - logical

OpenStudy (anonymous):

whatever u find easiest - thres really no 'right way'

OpenStudy (amistre64):

yep; but there are plenty of wrong ways lol

OpenStudy (anonymous):

yea